(60) 2. The Nd:YAG laser is used in eye surgery called posteriour capsulotomy, following cataract surgery, to correct cloudiness that affects vision. The laser emits light with a wavelength of 1064 nanometers (nm) in the infrared region of the electro-magnetic spectrum. Each laser pulse has 40.5 J of energy, contains many photons, and the pulse lasts for 30.5 nanoseconds (ns).

(15) 2a. Determine the frequency of the 1064 nm Nd:YAG light.

(15) 2b. Determine the energy of each photon (J/photon) of the 1064 nm Nd:YAG light

(15) 2c. Each laser pulse has 40.5 J of energy (40.5 J/pulse). Determine the number of photons in each laser pulse.

(15) 2d. The surgery requires 655 laser pulses. Determine the total number of moles of photons used in the surgery.
Answer #1

(2a) Frequency =(speed of the light)/wave length =(3*108)/(1064*10-9)=2.81 * 1014 Hz

(2b) Energy of each proton =plank constant *frequency =6.626*10-34 *2.81*1014 =1.86 * 10-19 J/Photon

(2c) Number of photons in laser pulse = 40.5/(1.86*10-19)= 2.177 *1020 photons.

(2d)Now 655 laser pulses contain 655*40.5= 26527.5 J of energy. So number of photon will be 26527.5/(1.86*10-19)= 1.426 *1023 photons.